OpenStudy (anonymous):

Solve for w. w^2-5w-36=0 enter answer in this format __,__

OpenStudy (amistre64):

the "format" there doesnt convey any meaningful information; that is simply the authors construction which is undefined outside of his parameters

OpenStudy (amistre64):

but given that; we need to factor the quadratic equation to see it easier

OpenStudy (amistre64):

x^2 + (b+c)x + (bc) is the format for this equation; so we need to find 2 numbers ; b and c that make this true: bc = -36 b+c = -5

OpenStudy (amistre64):

I can see from experience that -9 and +4 make this true -9(4) = -36 -9+4 = -5

OpenStudy (anonymous):

i thought i just have to figure out what (w)is

OpenStudy (amistre64):

so lets split this into: (x+b) (x+c) = 0 (x-9) (x+4) = 0 Another thing to keep in mind is that anything times 0 = 0; so there are 2 ways to get this to = 0; when x=9 we have: (9-9)(9+4)=0 and that is true when x=-4 we have: (-4-9)(-4+4) = 0 which is true

OpenStudy (amistre64):

yes, and this is how you do it .... but I used 'x' instead of 'w' just by instinct

OpenStudy (anonymous):

so 9,-4

OpenStudy (amistre64):

so; w = x in this .... its just a name, not a value. x=9 and x=-4 are your values yes, but I have no idea what the 'format' means so ill leave that for you to decipher ;)

OpenStudy (anonymous):

thank you (:

OpenStudy (amistre64):

youre welcome ☺